OpenStudy (anonymous):

the value of an uncirculated mint state 65 1950 jefferson nickel minted in denver is 3/2 the value a 1945 nickel minted in philadelphia is similar condition. together the total value of the two coins is $20. what is the value of each coin?

OpenStudy (ranga):

What is that 65 in the first line? 65 coins?

OpenStudy (anonymous):

it just says 'mint state 65' like that.

OpenStudy (ranga):

Assume the Philadelphia nickel is valued at "P" dollars and the Denver nickel is valued at "D" dollars. Denver nickel is 3/2 times the value of Philadelphia nickel. D = 3/2 * P ---- (1) Together the two nickels are worth $20. D + P = 20 ----- (2) Substitute D = 3/2P from (1) into (2) and solve for P. Then put P value in (1) and solve for D.

OpenStudy (anonymous):

so D+p=3/2 right?

OpenStudy (ranga):

Put D = 3/2P in (2): 3/2P + P = 20 P(3/2 + 1) = 20 P( (3+2) / 2 ) = 20 P * 5/2 = 20 P = 20 * 2 / 5 = 8 Put P = 8 in (1): D = 3/2 * P = 3/2 * 8 = 3 * 4 = 12 Philadelphia nickel is worth $8 Denver nickel is worth $12